OpenStudy (anonymous):

which of the following has a graph that passes through a point (3,1) a. 2x=y-7 b.y=4x-9 c.y=2/3x-5 d.3x-7y=2

OpenStudy (anonymous):

put 3 , 1 in the equation n find which one has LHS = RHS d gives Left hand side = right hand side when x = 3 and y = 1

OpenStudy (anonymous):

what is LHS = RHS? @chandanjha

OpenStudy (anonymous):

left hand side is LHS and right hand side is RHS 3x-7y = 3(3)- 7(1)=9-7=2 hence LHS=2 also RHS=2 hence d